Consider the country of Agrovia, whose GDP is discussed in “A Numerical Example” on page 396. Construct a set of national accounts like that in Table 20-6 assuming that wheat costs $5 per bushel, there is no depreciation, wages are three-fourths of national output, indirect business taxes are used to finance 100 percent of government spending, and the balance of income goes as rent income to farmers


1
Expert's answer
2021-08-30T12:35:46-0400

Output = $1000

Cost of FP = $750 in wages, $125 in interest, and $75 in rent


Profit  = output – expenses = $ 1000 – ($ 750 + $ 125 + $ 75) = $ 50

Investment & Consumption

"C = \u00be x 1000 = 750"

"I = 1000 \u2013 750 = 250"

Product Approaches   the GDP with Product Approaches is 1000

Income Approaches  "GDP\/Y = w + r + I + p"

"GDP\/Y= 750 + 125 + 75 + 50 = 1000"